OpenStudy (anonymous):

A truck can be rented from company A for $100 a day plus 0.30 per mile. Company B charges $80 a day plus 0.70 per mile to rent the same truck. How many miles must be driven to make the rental cost for Company A a better deal than Company B?

OpenStudy (anonymous):

@ConDawg @reemii @tkhunny @texaschic101

OpenStudy (anonymous):

turn into lieanr equations and set equal to eachoter

OpenStudy (anonymous):

\[100+0.30x=80+0.70x\]

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

We could graph two equations, one for each rental company, and compare where A is below B. A bit of a shortcut would be to set the two equal to each other, and find the amount of miles where they cost the same - 100+.3x = 80 + .7x 20 + .3x = .7x 20 = .4x 50 = x So, because B charges more per mile, after 50 miles, B costs more money, so A is a better deal.

OpenStudy (anonymous):

\[20+0.30x=0.70x\]

OpenStudy (anonymous):

\[20=0.4x\]

OpenStudy (anonymous):

\[x=50\]

OpenStudy (texaschic101):

x = number of miles 100 + .30x = 80 + .70x 100 - 80 = .70x - .30x 20 = .40x 20/.40 = x 50 = x Okay....they are the same at 50 mph. For company A to be a better deal, you would have to drive 51 miles....lets check Company A. 100 + .30(x) = 100 + .30(51) = 100 + 15.30 = 115.30 Company B 80 + .70(51) = 80 + 35.70 = 115.70 ANSWER : 51or more miles must be driven for Company A to be a better deal